# Context-Aware Weakly Supervised Image Manipulation Localization with SAM Refinement

### v1

@@ I Introduction
With the proliferation of maliciously forged images, the localization of manipulated regions has become critical for combating misinformation. Currently, most image manipulation localization methods leverage both image-level binary classification labels and pixel-level mask annotations for fully supervised training[1]. Early approaches leveraged noise, boundaries, and color inconsistencies [2, 3], while recent methods focus on content-agnostic features: CAT-Net [4] learns compression artifacts; MVSS-Net [5] uses multi-scale noise; Mantra-Net [6] detects operational boundaries; and PROMPT-IML [7] aligns semantic and high-frequency features via pre-trained models.
@@
Despite significant advances in fully supervised image manipulation localization methods[1], the growing diversity and complexity of manipulation techniques have made acquiring comprehensive pixel-level mask annotations increasingly impractical. Dense annotation of manipulated regions is both challenging and time-consuming. As a result, recent research has shifted towards weakly supervised approaches to enhance generalization and localization performance. For instance, WSCL[11] introduces multi-source consistency and inter-patch consistency modules by exploiting the self-consistency properties of manipulated images. While effective for manipulation detection tasks, its localization performance is limited by inadequate consideration of boundary information, which is crucial for accurate manipulation localization.
Image manipulation techniques like copy-move[12, 13] and splicing[14] involve non-generative operations in which manipulated regions are copied from authentic areas. As a result, the noise and texture within manipulated regions often resemble those of real regions, making the boundaries of manipulated areas crucial for localization. To better learn boundary information, we propose a Context-Aware Boundary Localization (CABL) module that utilizes context-inconsistency at the boundaries to more precisely identify manipulated regions. Furthermore, previous methods[11] in weakly supervised image manipulation localization have largely overlooked the potential of Class Activation Mapping (CAM)[15]. In our approach, we compute CAM maps using the output of a dual-branch Transformer-CNN network to generate coarse masks of manipulated regions. To refine these coarse masks, we leverage the strong capabilities of the Segment Anything Model (SAM)[16] through a CAM-Guided SAM Refinement (CGSR) module. Specifically, we use the bounding rectangle of the largest connected component from the coarse masks and the positive and negative sampling points from the CAM[15] without maximum-minimum normalization as visual prompts for SAM[16]. SAM[16] can generate more accurate localization masks using these visual prompts, without requiring additional training. By integrating these two modules, we propose a novel framework that combines CABL and CGSR within a dual-branch Transformer-CNN architecture. This framework can learn the context-inconsistency of the manipulated images under weakly supervised training conditions, enabling precise localization of the manipulated regions.

### v2

@@ I Introduction
With the proliferation of maliciously forged images, the localization of manipulated regions has become critical for combating misinformation. Currently, most image manipulation localization methods leverage both image-level binary classification labels and pixel-level mask annotations for fully supervised training[1, 2]. Early approaches leveraged noise, boundaries, and color inconsistencies [3, 4], while recent methods focus on content-agnostic features: CAT-Net [5] learns compression artifacts; MVSS-Net [6] uses multi-scale noise; Mantra-Net [7] detects operational boundaries; and PROMPT-IML [8] aligns semantic and high-frequency features via pre-trained models.
@@
Despite significant advances in fully supervised image manipulation localization methods[1, 12], the growing diversity and complexity of manipulation techniques have made acquiring comprehensive pixel-level mask annotations increasingly impractical. Dense annotation of manipulated regions is both challenging and time-consuming. As a result, recent research has shifted towards weakly supervised approaches to enhance generalization and localization performance. For instance, WSCL[13] introduces multi-source consistency and inter-patch consistency modules by exploiting the self-consistency properties of manipulated images. While effective for manipulation detection tasks, its localization performance is limited by inadequate consideration of boundary information, which is crucial for accurate manipulation localization.
Image manipulation techniques like copy-move[14, 15] and splicing[16] involve non-generative operations in which manipulated regions are copied from authentic areas. As a result, the noise and texture within manipulated regions often resemble those of real regions, making the boundaries of manipulated areas crucial for localization. To better learn boundary information, we propose a Context-Aware Boundary Localization (CABL) module that utilizes context-inconsistency at the boundaries to more precisely identify manipulated regions. Furthermore, previous methods[13, 17] in weakly supervised image manipulation localization have largely overlooked the potential of Class Activation Mapping (CAM)[18]. In our approach, we compute CAM maps using the output of a dual-branch Transformer-CNN network to generate coarse masks of manipulated regions. To refine these coarse masks, we leverage the strong capabilities of the Segment Anything Model (SAM)[19] through a CAM-Guided SAM Refinement (CGSR) module. Specifically, we use the bounding rectangle of the largest connected component from the coarse masks and the positive and negative sampling points from the CAM[18] without maximum-minimum normalization as visual prompts for SAM[19]. SAM[19] can generate more accurate localization masks using these visual prompts, without requiring additional training. By integrating these two modules, we propose a novel framework that combines CABL and CGSR within a dual-branch Transformer-CNN architecture. This framework can learn the context-inconsistency of the manipulated images under weakly supervised training conditions, enabling precise localization of the manipulated regions.